# Duality theory in linear optimization and its extensions — formally verified

### v1

@@
  Abstract:  Farkas established that a system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. We state and formally prove several Farkas-like theorems over linearly ordered fields in Lean 4. Furthermore, we extend duality theory to the case when some coefficients are allowed to take “infinite values”. Keywords:  Farkas lemma, linear programming, extended rationals, extended reals, calculus of inductive constructions
@@ 1 Introduction
When studying linear programming, we often ask whether a given system of linear inequalities has a solution. Duality theorems answer these naturally arising questions as follows: A system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. When we formulate duality theorems as “either there exists a solution, or there exists a vector of coefficients that tells us how to derive $0<0$ from given inequalities”, they are usually called theorems of alternatives. Two well-known theorems of alternatives are as follows (Farkas [9, 10]; Minkowski [18]).
@@
  • We establish (and formally prove in Lean 4) a new generalization of LABEL:inequalityFarkas and LABEL:StandardLP.strongDuality to the case when some of the coefficients are allowed to have infinite values. This scenario can be motivated by discrete optimization problems with “hard” constraints (the “hard” constraints declare what has to be be satisfied, whereäs “soft” constraints declare what should be optimized). A common way to concisely write down such problems mathematically is to use infinite coefficients in front of the corresponding terms in the objective. Since infinities are not allowed in traditional LPs, “soft” and “hard” constraints need to be handled differently when formulating LP relaxations of such problems (see e.g. [13]). Our work provides a more direct way to formulate such relaxations. Section 1.2 says more about the extensions.
@@ 1.1 Bartl’s generalization
Let $J$ be a finite type. Let $R$ be a linearly ordered division ring. Let $W$ be an $R$-module. Let $A$ be an $R$-linear map from from $W$ to $(J\rightarrow R)$. Let $b$ be an $R$-linear map from from $W$ to $R$. Exactly one of the following exists:
@@
In the last branch, $A~{}y\geq 0$ uses the partial order on $(J\rightarrow R)$ whereäs $b~{}y<0$ uses the linear order on $V$. Note that LABEL:fintypeFarkasBartl subsumes LABEL:scalarFarkas (as well as the other versions based on equality), since $R$ can be viewed as a linearly ordered module over itself. We prove LABEL:fintypeFarkasBartl in Section 4, which is where the heavy lifting comes. Our proof is based on [3].
@@ 00     $\bot$ 00 00
It is straightforward to check that the dual of a valid LP is valid.
@@ 1.3 Structure of this paper
Repository https://github.com/madvorak/duality/tree/v2 contains the full version of all definitions, statements, and proofs. They are written in a formal language called Lean, which provides a guarantee that every step of every proof follows from valid logical axioms.111 The only axioms used in our proofs are propext, Classical.choice, and Quot.sound, which you can check by the #print axioms command. We use Lean version 4.11.0 together with the Lean mathematical library Mathlib [17] revision 00df099 (dated 2024-09-02). This paper attempts to be an accurate description of the duality project. However, in case of any discrepancy, the code shall prevail.
@@ 2 Preliminaries
There are many layers of definitions that are built before say what a linearly ordered field $F$ is (used e.g. in LABEL:equalityFarkas), what a linearly ordered division ring $R$ is (used e.g. in LABEL:scalarFarkas), and what a linearly ordered abelian group $V$ is (used in LABEL:fintypeFarkasBartl). Section 2.1 mainly documents existing Mathlib definitions; we also highlight how we added linearly ordered division rings into our project. Section 2.2 then explains how we defined extended linearly ordered fields. Section 2.3 reviews vectors and matrices; we especially focus on how multiplication between them is defined. Section 2.4 explains how modules are implemented. Section 2.5 provides the formal definition of extended linear programs.
@@ 2.1 Review of algebraic typeclasses that our project depends on
Additive semigroup is a structure on any type with addition where the addition is associative:
@@
Similarly, semigroup is a structure on any type with multiplication where the multiplication is associative:
@@
Additive monoid is an additive semigroup with the “zero” element that is neutral with respect to addition from both left and right, thanks to which we can define a scalar multiplication by the natural numbers:
@@
Similarly, monoid is a semigroup with the “one” element that is neutral with respect to multiplication from both left and right, thanks to which we can define a power to the natural numbers:
@@
zsmul_succ (n : $\mathbb{N}$) (a : G) : zsmul (Int.ofNat n.succ) a = zsmul (Int.ofNat n) a + a
@@
zpow_succ (n : $\mathbb{N}$) (a : G) : zpow (Int.ofNat n.succ) a = zpow (Int.ofNat n) a * a
@@
Distrib is a structure on any type with addition and multiplication where both left distributivity and right distributivity hold:
@@
Additive monoid with one and abelian monoid with one are defined from additive monoid and abelian monoid, equipped with the symbol “one” and embedding of natural numbers (please ignore the difference between Type u and Type* as it is the same thing for all our purposes):
@@
class DivisionRing ($\alpha$ : Type*) extends Ring $\alpha$, DivInvMonoid $\alpha$, Nontrivial $\alpha$, NNRatCast $\alpha$, RatCast $\alpha$ where
mul_inv_cancel : $\forall$ (a : $\alpha$), a $\neq$ 0 $\rightarrow$ a * a-1 = 1
inv_zero : (0 : $\alpha$)-1 = 0
@@
Partial order is a reflexive & antisymmetric & transitive relation:
@@ 2.3 Vectors and matrices
instance Pi.hasLe $\iota$ : Type u} $\pi$ : $\iota$rightarrow Type v} [$\forall$ i, LE ($\pi$ i)] :
@@ 2.3.1 Mathlib definitions versus our definitions
def Matrix.dotProduct [Fintype m] [Mul $\alpha$] [AddCommMonoid $\alpha$] (v w : m $\rightarrow$ $\alpha$) : $\alpha$ :=
@@
Infix notation for these two operations is defined as follows (the keyword infixl when defining an operator $\circ$ means that x $\circ$ y $\circ$ z gets parsed as (x $\circ$ y) $\circ$ z whether or not it makes sense, whereäs the keyword infixr when defining an operator $\circ$ means that A $\circ$ B $\circ$ v gets parsed as A $\circ$ (B $\circ$ v) whether or not it makes sense; the numbers 72 and 73 determine precedence):
@@
def Matrix.dotProd (v : I $\rightarrow$ $\alpha$) (w : I $\rightarrow$ $\gamma$) : $\alpha$ :=
@@
infixl:72 " $\cdot$ " => Matrix.dotProd
@@
In the infix notation, you can distinguish between the standard Mathlib definitions and our definitions by observing that Mathlib operators put the letter v to the right of the symbol whereäs our operators put a letter to the left of the symbol.
We find it unfortunate that the Mathlib name of dotProduct is prefixed by the Matrix namespace. While Matrix.mulVec allows for the use of dot notation in places where infix notation is not used, the full name Matrix.dotProduct only clutters the code. Even though we do not like it, we decided to namespace our definitions in the same way for consistency.
Since we have new definitions, we have to rebuild all API (a lot of lemmas) for Matrix.dotProd and Matrix.mulWeig from scratch. This process was very tiresome. We decided not to develop a full reusable library, but prove only those lemmas we wanted to use in our project. For similar reasons, we did not generalize the Mathlib definition of “vector times matrix”, as “matrix times vector” was all we needed. It was still a lot of lemmas.
@@ 2.4 Modules and how to order them
Given types $\alpha$ and $\beta$ such that $\alpha$ has a scalar action on $\beta$ and $\alpha$ forms a monoid, Mathlib defines multiplicative action where $1$ of type $\alpha$ gives the identity action and multiplication in the monoid associates with the scalar action:
@@ 2.5 Linear programming
def ExtendedLP.IsBoundedBy (P : ExtendedLP I J F) (r : F) : Prop :=
@@
def ExtendedLP.IsUnbounded (P : ExtendedLP I J F) : Prop :=
@@
3 Formal statements of selected results
@@ 3.2 Main results
theorem ExtendedLP.strongDuality (P : ExtendedLP I J F) (hP : P.IsFeasible $\vee$ P.dualize.IsFeasible) :
@@ 6 Proving the Extended strong LP duality
Lemma (no_bot_of_feasible ).
Let $F$ be a linearly ordered field. Let $P$ be a valid linear program over $F_{\infty}$. If $P$ is feasible, then $b$ does not contain $\bot$.
Proof idea: This is just a weaker version of LABEL:no-bot-of-reaches.
@@
Lemma LABEL:no-bot-of-feasible is used 12 times throughout this proof.
@@
Lemma (strongDuality_of_prim_feas ).
@@
Lemma (strongDuality_of_dual_feas ).
@@ 7.1 Counterexamples for LABEL:extendedFarkas
We also claimed in Section 1.2 that changing condition $(-A^{T})\cdot y\leq 0$ to $A^{T}\cdot y\geq 0$ in LABEL:extendedFarkas would cause the theorem to fail even when $A$ has a single $\bot$ entry. The counterexample is as follows:
@@ 8.2 Hahn-Banach theorems
Several Hahn-Banach theorems have been formalized in Lean as a part of Mathlib. In particular, it would have been be possible to prove LABEL:equalityFarkas for reals using the Hahn-Banach separation theorem for a convex closed set and a point [5]. It would then have been possible to extend the result to other duality theorems for reals. To our knowledge, the theorem [5] cannot be used to prove LABEL:equalityFarkas for an arbitrary linearly ordered field. Therefore, we decided to prove LABEL:equalityFarkas without appealing to the geometry and topology.

### v2

@@
  Abstract:  Farkas established that a system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. We state and formally prove several Farkas-like theorems over linearly ordered fields in Lean 4. Furthermore, we extend duality theory to the case when some coefficients are allowed to take “infinite values”. Keywords:  Farkas lemma, linear programming, extended reals, calculus of inductive constructions
@@ 1 Introduction
In optimization and related fields, we often ask whether a given system of linear inequalities has a solution. Duality theorems answer these naturally arising questions as follows: A system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. When we formulate duality theorems as “either there exists a solution, or there exists a vector of coefficients that tells us how to derive $0<0$ from given inequalities”, they are usually called theorems of alternatives. Two well-known theorems of alternatives are as follows (Farkas [12, 13]; Minkowski [21]).
@@
  • We establish (and formally prove in Lean 4) a new generalization of LABEL:inequalityFarkas and LABEL:StandardLP.strongDuality to the case when some of the coefficients are allowed to have infinite values. This scenario can be motivated by discrete optimization problems with “hard” constraints (the “hard” constraints declare what has to be be satisfied, whereas “soft” constraints declare what should be optimized). A common way to concisely write down such problems mathematically is to use infinite coefficients in front of the corresponding terms in the objective. Since infinities are not allowed in traditional LPs, “soft” and “hard” constraints need to be handled differently when formulating LP relaxations of such problems (see e.g. [16]). Our work provides a more direct way to formulate such relaxations. Section 1.2 says more about the extensions.
@@ 1.1 Bartl’s generalization
Let $J$ be a finite type. Let $R$ be a linearly ordered division ring. Let $W$ be an $R$-module. Let $A$ be an $R$-linear map from $W$ to $(J\rightarrow R)$. Let $b$ be an $R$-linear map from $W$ to $R$. Exactly one of the following exists:
@@
In the last branch, $A~{}y\geq 0$ uses the partial order on $(J\rightarrow R)$ whereas $b~{}y<0$ uses the linear order on $V$. Note that LABEL:fintypeFarkasBartl subsumes LABEL:scalarFarkas (as well as the other versions based on equality), since $R$ can be viewed as a linearly ordered module over itself. We prove LABEL:fintypeFarkasBartl in Section 4, which is where the heavy lifting comes. Our proof is based on [5].
@@ 00     $\bot$ 00 00
It is straightforward to check that the dual of a valid LP is valid. Later we show that a valid $P$ cannot have $P^{\star}$ undefined.
@@
1.2.1 Example — cheap lunch
According to [1] a kilogram of boiled white rice contains $27$ g protein and $1300$ kcal. According to [2] a kilogram of boiled lentils contains $90$ g protein and $1150$ kcal. Let’s say that a kilogram of boiled white rice costs $0.92$ euro and that a kilogram of boiled lentils costs $1.75$ euro. We want to cook a lunch, as cheap as possible, that contains at least $30$ g protein and at least $700$ kcal. The choice of white rice and lentils isn’t random — one of the authors ate this lunch at a mathematical camp — at the moment, he didn’t like the lunch — it wasn’t very tasty — but later he realized what an awesome nutritional value the lunch had for its low price. This is a simple example of the well-known diet problem [11].
$\begin{array}[]{rcl}\min&&0.92\cdot r+1.75\cdot l\\ (-27)\cdot r+(-90)\cdot l&\leq&-30\\ (-1300)\cdot r+(-1150)\cdot l&\leq&-700\\ r&\geq&0\\ l&\geq&0\end{array}$
Using a numerical LP solver, we obtain the optimal solution $r=0.331588$ and $l=0.233857$ which satisfies our dietary requirements for $0.714311$ euro. The dual of this problem, in the sense of LABEL:ExtendedLP.strongDuality, is the following LP:
$\begin{array}[]{rcl}\min&&(-30)\cdot p+(-700)\cdot k\\ 27\cdot p+1300\cdot k&\leq&0.92\\ 90\cdot p+1150\cdot k&\leq&1.75\\ p&\geq&0\\ k&\geq&0\end{array}$
Using a numerical LP solver, we obtain the optimal solution $p=0.0141594$ and $k=0.000413613$ leading to the objective value $-0.714311$ here. To summarize, the cheapest lunch that contains at least 30 g protein and at least 700 kcal consists of $332$ g white rice and $224$ g lentils; the shadow cost of a gram of protein is $0.014$ euro and the shadow cost of a kcal is $0.00041$ euro in our settings.
Now consider settings where lentils are out of stock. We still want to obtain at least 30 g protein and at least 700 kcal. What is the price of our lunch now? On paper, it is most natural to entirely remove lentils out of the picture and recompute the LP. However, computer proof systems generally don’t like it when sizes of matrices change, so let’s do the modification of input in place. One could suggest that, in order to model the newly arised situation, the price of lentils will be increased to $999999$ euro, so that the optimal solution will assign 00 to it. This is, however, not mathematically elegant, as it requires some engineering insight into setting the constant to an appropriately large number. We claim that the proper mathematical approach is to increase the price of lentils to infinity! Let’s see how our original LP will look now:
$\begin{array}[]{rcl}\min&&0.92\cdot r+\top\cdot l\\ (-27)\cdot r+(-90)\cdot l&\leq&-30\\ (-1300)\cdot r+(-1150)\cdot l&\leq&-700\\ r&\geq&0\\ l&\geq&0\end{array}$
The optimal solution is now $r=1.111111$ and $l=0$ and costs $1.022222$ euro. The dual of this problem is the following LP:
$\begin{array}[]{rcl}\min&&(-30)\cdot p+(-700)\cdot k\\ 27\cdot p+1300\cdot k&\leq&0.92\\ 90\cdot p+1150\cdot k&\leq&\top\\ p&\geq&0\\ k&\geq&0\end{array}$
At this point, the shadow price of a gram of protein is $0.034074$ euro but the shadow price of kcal is 00 euro. The objective value is $-1.022222$ in accordance with our theoretical prediction.
@@ 1.3 Structure of this paper
Repository https://github.com/madvorak/duality/tree/v3.2 contains the full version of all definitions, statements, and proofs. They are written in a formal language called Lean, which provides a guarantee that every step of every proof follows from valid logical axioms.111 The only axioms used in our proofs are propext, Classical.choice, and Quot.sound, which you can check by the #print axioms command. We use Lean version 4.18.0 together with the Lean mathematical library Mathlib [20] revision aa936c3 (dated 2025-04-01). This paper attempts to be an accurate description of the duality project. However, in case of any discrepancy, the code shall prevail.
@@ 2 Preliminaries
There are many layers of definitions that are built before say what a linearly ordered field is (used e.g. in LABEL:equalityFarkas), what a linearly ordered division ring is (used e.g. in LABEL:scalarFarkas), and what a linearly ordered abelian group is (used in LABEL:fintypeFarkasBartl). Section 2.1 mainly documents existing Mathlib definitions (but omitting declarations of default values); we also highlight how we added linearly ordered division rings into our project. Section 2.2 then explains how we defined extended linearly ordered fields. Section 2.3 reviews vectors and matrices; we especially focus on how multiplication between them is defined. Section 2.4 explains how modules are implemented. Section 2.5 provides the formal definition of extended linear programs.
@@ 2.1 Review of algebraic typeclasses that our project depends on
Additive semigroup is a structure on any type with addition (denoted by the infix + operator) where the addition is associative:
@@
Semigroup is a structure on any type with multiplication (denoted by the infix * operator) where the multiplication is associative:
@@
Additive monoid is an additive semigroup with the “zero” element that is neutral with respect to addition from both left and right, equipped with a scalar multiplication by the natural numbers:
@@
Similarly, monoid is a semigroup with the “one” element that is neutral with respect to multiplication from both left and right, equipped with a power to the natural numbers:
@@
zsmul_succ (n : $\mathbb{N}$) (a : G) : zsmul n.succ a = zsmul n a + a
@@
zpow_succ (n : $\mathbb{N}$) (a : G) : zpow n.succ a = zpow n a * a
@@
Distrib is a structure on any type with addition and multiplication where both left distributivity and right distributivity hold (please ignore the difference between Type u and Type* as it is the same thing for all our purposes):
@@
Additive monoid with one and abelian monoid with one are defined from additive monoid and abelian monoid, equipped with the symbol “one” and embedding of natural numbers:
@@
class DivisionRing (K : Type*) extends Ring K, DivInvMonoid K, Nontrivial K, NNRatCast K, RatCast K where
mul_inv_cancel : $\forall$ (a : K), a $\neq$ 0 $\rightarrow$ a * a-1 = 1
inv_zero : (0 : K)-1 = 0
@@
Partial order is an antisymmetric preoder (that is, a reflexive & antisymmetric & transitive relation):
@@ 2.3 Vectors and matrices
instance Pi.hasLe $\iota$ : Type*} $\pi$ : $\iota$rightarrow Type*} [$\forall$ i, LE ($\pi$ i)] :
@@ 2.3.1 Mathlib definitions versus our definitions
def dotProduct [Fintype m] [Mul $\alpha$] [AddCommMonoid $\alpha$] (v w : m $\rightarrow$ $\alpha$) : $\alpha$ :=
@@
Infix notation for these two operations is defined as follows (the keyword infixl when defining an operator $\circ$ means that x $\circ$ y $\circ$ z gets parsed as (x $\circ$ y) $\circ$ z whether or not it makes sense, whereas the keyword infixr when defining an operator $\circ$ means that A $\circ$ B $\circ$ v gets parsed as A $\circ$ (B $\circ$ v) whether or not it makes sense; the numbers 72 and 73 determine precedence — higher number means tighter):
@@
def dotWeig (v : I $\rightarrow$ $\alpha$) (w : I $\rightarrow$ $\gamma$) : $\alpha$ :=
@@
infixl:72 " $\cdot$ " => dotWeig
@@
In the infix notation, you can distinguish between the standard Mathlib definitions and our definitions by observing that Mathlib operators put the letter v to the right of the symbol whereas our operators put a letter to the left of the symbol.
Since we have new definitions, we have to rebuild all API (a lot of lemmas) for dotWeig and Matrix.mulWeig from scratch. This process was very tiresome. We decided not to develop a full reusable library, but prove only those lemmas we wanted to use in our project. For similar reasons, we did not generalize the Mathlib definition of “vector times matrix”, as “matrix times vector” was all we needed. It was still a lot of lemmas.
@@ 2.4 Modules and how to order them
Given types $\alpha$ and $\beta$ such that $\alpha$ has a scalar action on $\beta$ (denoted by the infix • operator) and $\alpha$ forms a monoid, Mathlib defines multiplicative action where $1$ of type $\alpha$ gives the identity action and multiplication in the monoid associates with the scalar action:
@@ 2.5 Linear programming
def ExtendedLP.IsBoundedBy [Fintype J] (P : ExtendedLP I J F) (r : F) : Prop :=
@@
def ExtendedLP.IsUnbounded [Fintype J] (P : ExtendedLP I J F) : Prop :=
@@
3 Formal statements of selected theorems
@@ 3.2 Main results
theorem ValidELP.strongDuality (P : ValidELP I J F) (hP : P.IsFeasible $\vee$ P.dualize.IsFeasible) :
@@ 6 Proving the Extended strong LP duality
Lemma (strongDuality_of_prim_feasible ).
@@
Lemma (strongDuality_of_dual_feasible ).
@@ 7.1 Counterexamples for LABEL:extendedFarkas
We also claimed in Section 1.2 that changing condition $(-A^{T})\cdot y\leq 0$ to $A^{T}\cdot y\geq 0$ in LABEL:extendedFarkas would cause the theorem to fail even when $A$ has only a single $\bot$ entry. The counterexample is as follows:
@@ 8.2 Hahn-Banach theorems
Several Hahn-Banach theorems have been formalized in Lean as a part of Mathlib. In particular, it would have been possible to prove LABEL:equalityFarkas for reals using the Hahn-Banach separation theorem for a convex closed set and a point [7]. It would then have been possible to extend the result to other duality theorems for reals. To our knowledge, the theorem [7] cannot be used to prove LABEL:equalityFarkas for an arbitrary linearly ordered field. Therefore, we decided to prove LABEL:equalityFarkas without appealing to the geometry and topology.